OpenStudy (anonymous):

Find the prime factorization of the number. 144

OpenStudy (anonymous):

2*2*2*2*3*3

OpenStudy (anonymous):

Or \[\Large 2{^4}\times 3{^2}\]

OpenStudy (jamesj):

Exactly. You know that \(144 = 12^2\) and \(12 = 4 \times 3 = 2^2 3\). Hence \( 144 = (2^2 3)^2 = 2^{2\times2} 3^2 = 2^43^2 \)